OpenStudy (anonymous):

The slope of a line is -1/3, and the y-intercept is 10/3. What is the equation of the line written in general form? 10x + 3y - 1 = 0 x + 3y + 10 = 0 x + 3y - 10 = 0 i got as far as y=-1/3x+10/3

OpenStudy (anonymous):

where do i go from there

OpenStudy (anonymous):

@texaschic101

OpenStudy (texaschic101):

y = -1/3x + 10/3 multiply the equation by 3 and tell me what you get

OpenStudy (texaschic101):

by doing this, it will get rid of the fractions

OpenStudy (anonymous):

kk brb

OpenStudy (anonymous):

i got y=-1x+10

OpenStudy (texaschic101):

you have to multiply the y by 3 as well

OpenStudy (anonymous):

so 3=-1x+10 ?

OpenStudy (anonymous):

3y*

OpenStudy (texaschic101):

3y = -x + 10 now add x to both sides and subtract 10 from both sides and set it equal to zero

OpenStudy (texaschic101):

we will do it in steps.....add x to both sides. This will cancel out the x's on the right side

OpenStudy (anonymous):

can you just show me how to do it im having trouble

OpenStudy (texaschic101):

3y = -x + 10 (add x to both sides) 3y + x = -x + x + 10 combine like terms 3y + x = 10 now subtract 10 from both sides 3y + x - 10 = 10 - 10 combine like terms 3y + x -10 = 0 x has to be first so change the equation around... x + 3y - 10 = 0

OpenStudy (texaschic101):

understand ?

OpenStudy (anonymous):

yeh im writing all the steps down rite now

OpenStudy (texaschic101):

if you have any questions, feel free to ask :)

OpenStudy (anonymous):

kk thanks for helping

OpenStudy (texaschic101):

no problem :)